# HyperAIRI: a plug-and-play algorithm for precise hyperspectral image reconstruction in radio interferometry

### v1

@@ Abstract
The next-generation radio-interferometric (RI) telescopes require imaging algorithms capable of forming high-resolution high-dynamic-range images from large data volumes spanning wide frequency bands. Recently, AIRI, a plug-and-play (PnP) approach taking the forward-backward algorithmic structure (FB), has demonstrated state-of-the-art performance in monochromatic RI imaging by alternating a data-fidelity step with a regularisation step via learned denoisers. In this work, we introduce HyperAIRI, its hyperspectral extension, underpinned by learned hyperspectral denoisers enforcing a power-law spectral model. For each spectral channel, the HyperAIRI denoiser takes as input its current image estimate, alongside estimates of its two immediate neighbouring channels and the spectral index map, and provides as output its associated denoised image. To ensure convergence of HyperAIRI, the denoisers are trained with a Jacobian regularisation enforcing non-expansiveness. To accommodate varying dynamic ranges, we assemble a shelf of pre-trained denoisers, each tailored to a specific dynamic range. At each HyperAIRI iteration, the spectral channels of the target image cube are updated in parallel using dynamic-range-matched denoisers from the pre-trained shelf. The denoisers are also endowed with a spatial image faceting functionality, enabling scalability to varied image sizes. Additionally, we formally introduce Hyper-uSARA, a variant of the optimisation-based algorithm HyperSARA, promoting joint sparsity across spectral channels via the $\ell_{2,1}$-norm, also adopting FB. We evaluate HyperAIRI’s performance on simulated and real observations. We showcase its superior performance compared to its optimisation-based counterpart Hyper-uSARA, CLEAN’s hyperspectral variant in WSClean, and the monochromatic imaging algorithms AIRI and uSARA. HyperAIRI’s MATLAB implementation is available in the BASPLib code library.
keywords:
techniques: image processing – techniques: interferometric.
pubyear: 2025pagerange: HyperAIRI: a plug-and-play algorithm for precise hyperspectral image reconstruction in radio interferometryB
1 Introduction
@@
Radio interferometry is a powerful technique in astronomy, providing high-resolution views of the sky by combining signals from an array of antennas. In particular, the upcoming Square Kilometre Array (SKA, Scaife, 2020) is designed to deliver unprecedented sensitivity and resolution, with phase 1 offering more than a tenfold improvement compared to its precursor instruments, MeerKAT (Jonas, 2009) and Australian SKA Pathfinder (ASKAP, Hotan et al., 2021). These telescopes observe across wide frequency bands by collecting Fourier measurements, enabling the study of astrophysical processes at both spatial and spectral scales.
@@
However, the associated inverse problem of hyperspectral image reconstruction is inherently ill-posed due to incomplete Fourier sampling, instrumental noise, and calibration errors. Monochromatic algorithms, which reconstruct spectral channels independently, achieve only limited recovery under such conditions. In contrast, joint-channel reconstruction exploits spectral correlations to regularise the ill-posed problem and enables high-precision recovery, particularly in high-dynamic-range (HDR) scenarios. Nevertheless, the sheer scale of hyperspectral data and the need for precise imaging significantly increase the computational demands, calling for efficient and scalable algorithms with joint-channel reconstruction capability.
@@
Most radio sources emit synchrotron radiation that can be described by a power-law spectral model (Thompson et al., 2017). Building on the celebrated CLEAN algorithm (Högbom, 1974), its first hyperspectral extension, Multi-Frequency (MF) CLEAN, models spectral variations via polynomial functions, and solves for the corresponding coefficient images during reconstruction (Sault & Wieringa, 1994). Further improvements are enabled by combining MF with Multi-Scale (MS) CLEAN (Cornwell, 2008), enhancing the recovery of extended emission (Rau & Cornwell, 2011; Offringa & Smirnov, 2017). However, low-order polynomials are often inadequate to capture complex spectral behaviour (Scaife & Heald, 2012). To address this, Ceccotti et al. (2023) have injected a forced-spectrum fitting approach within CLEAN, enforcing a prior spectral knowledge directly during reconstruction. While computationally efficient, CLEAN-based algorithms still fall short in achieving precise imaging, i.e. high-resolution and high-sensitivity imaging. These limitations stem from their simplistic prior model: spatial resolution remains constrained by the instrument’s nominal resolution, and dynamic range is degraded by the addition of residuals in the final reconstruction.
@@
Bayesian inference methods, such as the variational Bayes approach underpinning the RESOLVE algorithm (Junklewitz et al., 2016), have demonstrated superior precision and robustness over CLEAN-based techniques (Arras et al., 2021; Roth et al., 2023). Assuming a first-order power-law spectral model, Junklewitz et al. (2015) proposed the joint estimation of a reference-frequency image and a spectral index map. Although these methods naturally provide uncertainty quantification functionality, their scalability and precision on large-scale observations remain to be demonstrated.
@@
Methods rooted in optimisation theory have shown promise in astronomical imaging, offering precision imaging. Optimisation provides a robust and versatile framework to address ill-posed inverse problems, typically formulated as minimisation tasks solved via convergent iterative algorithmic structures. In this context, the reconstructed image is obtained as the minimiser of an objective function composed of a data fidelity function and a regularisation function encoding sophisticated image models. Optimisation offers a multitude of distributed algorithmic structures for parallelised reconstruction at scale. Among these methods, the SARA family of algorithms leverages sparsity-based priors, demonstrating superior imaging quality compared to CLEAN across diverse imaging scenarios (Carrillo et al., 2012; Onose et al., 2016, 2017; Repetti et al., 2017; Birdi et al., 2018; Pratley et al., 2018; Dabbech et al., 2018; Terris et al., 2023; Wilber et al., 2023a). Specifically, HyperSARA, promotes joint-channel average sparsity of the hyperspectral image cube in a redundant wavelet dictionary, and its low-rankness (Abdulaziz et al., 2016, 2019). HyperSARA takes the primal-dual forward-backward algorithmic structure (PD-FB, Condat, 2013), which enables full splitting of the functions involved, and parallel updates of the dual variables. To further improve scalability to large image sizes, Thouvenin et al. (2023a) proposed its faceted variant operating on 3D spatio-spectral image facets. The algorithm has been validated on real observational data, demonstrating state-of-the-art reconstruction (Thouvenin et al., 2023b). Yet, HyperSARA requires heavy cross-channel communications and could introduce significant overhead for extreme-scale problems. In this work, we introduce another simpler variant of HyperSARA, dubbed Hyper-uSARA, which adopts an unconstrained formulation of the minimisation task, thereby eliminating the need for exact noise distribution knowledge. It also takes a simpler algorithmic structure, the FB (Combettes & Pesquet, 2011), alternating data fidelity and regularisation steps, while still allowing for full parallelisation within each step.
@@
Recently, deep learning has gained increasing attention in RI imaging for its ability to produce reconstructions with rapid inference. Existing approaches focus on monochromatic imaging. These include traditional, end-to-end neural networks, either reconstructing images from back-projected data (Connor et al., 2022; Gheller & Vazza, 2022) or post-processing sub-optimal outputs from other algorithms such as CLEAN (Terris et al., 2019). However, these methods often lack precision, robustness, and interpretability (Goodfellow et al., 2014; Nguyen et al., 2015; Pang et al., 2018). Conditioned denoising diffusion probabilistic models have also been explored for RI imaging (Wang et al., 2023; Drozdova et al., 2024). They face similar issues and remain in early development.
@@
Advanced architectures such as unrolling Deep Neural Networks (DNN) promote data consistency and possibly algorithmic stability by unfolding a fixed number of iterations of an optimisation algorithm as key modules within the network (Monga et al., 2021). These networks are trained end-to-end while preserving an explicit link to optimisation theory. However, existing unrolling schemes remain limited by their fixed iteration count, and lack formal convergence guarantees. Furthermore, embedding computationally-intensive measurement operators directly into the network can make training prohibitively expensive.
@@
In contrast, the Plug-and-Play (PnP) framework bridges optimisation theory and deep learning (Venkatakrishnan et al., 2013; Romano et al., 2017; Reehorst & Schniter, 2018) offering a more flexible and robust alternative. In this approach, proximal operators enforcing handcrafted regularisations within an optimisation algorithm are replaced by learned DNN denoisers. These denoisers can be agnostic to the measurement process, making them adaptable to diverse configurations in RI imaging. Among these, the AIRI family of PnP algorithms (Terris et al., 2023, 2025) for monochromatic RI imaging, endowed with convergence guarantees (Pesquet et al., 2021). AIRI has been validated on both simulated and real wide-field observational data (Dabbech et al., 2022; Wilber et al., 2023b). Thanks to their convolutional architecture and compact receptive fields, AIRI denoisers are typically trained on small image patches, and seamlessly applied to images of arbitrary size during inference. Additionally, they can be easily applied in parallel to image facets for scalability. These properties make AIRI well suited to large-scale, precision imaging.
@@
Inspired by unrolled networks and PnP algorithms, Aghabiglou et al. (2024) have recently proposed R2D2, a novel imaging paradigm for radio interferometry. R2D2 is underpinned by a series of iteration-specific DNNs, each taking as input the current image estimate and associated data residual. The final reconstruction is formed by summing the DNNs’ output. In contrast to unrolled networks, R2D2 decouples data fidelity from the network architecture thereby avoiding the associated computational challenges. Conceptually, it can be interpreted as a learned version of CLEAN (Dabbech et al., 2024), substituting CLEAN’s minor cycles with iteration-specific DNNs. The novel paradigm has demonstrated superior imaging precision to the state of the art in field, namely AIRI and uSARA, both in simulation and on real data, only at a fraction of their computational cost. However, early incarnations of R2D2 have been so far trained in a telescope-specific setting, and for monochromatic intensity imaging at small scale. Extending it to large scale wideband imaging poses additional challenges that remain to be addressed.
@@
Building on the precision and scalability of AIRI, we propose HyperAIRI, a novel PnP algorithm tailored for high-precision hyperspectral RI imaging. HyperAIRI applies learned denoisers to each spectral channel in parallel with side information from neighbouring channels. The reconstruction approach further enhances spectral coherence by incorporating a physical power-law model, with the spectral index map either estimated on the fly during reconstruction or provided as an input. This design ensures local communication at each iteration and requires global communication only during spectral index updates, making HyperAIRI highly scalable and efficient in distributed high-performance computing (HPC) environments. Similarly to AIRI, HyperAIRI takes FB, with convergence guaranteed by the non-expansiveness of the denoisers, enforced via a Jacobian regularisation term in the training loss. We provide a shelf of pre-trained denoisers covering a wide dynamic-range, eliminating the need for problem-specific training. We demonstrate the superior performance of HyperAIRI, compared with state-of-the-art methods in the field, including the optimisation-based Hyper-uSARA and joint-channel CLEAN, on simulated the Very Large Array (VLA, Perley et al., 2011) data and real ASKAP observations.
@@
The remainder of this paper is organised as follows. In Section 2, we formalise the hyperspectral RI imaging problem. Section 3 reviews the hyperspectral SARA prior and formulates the Hyper-uSARA algorithm. Section 4 presents the HyperAIRI algorithm, including the network architecture, training methodology, and overall algorithm structure. Section 5 reports experimental results on simulated datasets to assess HyperAIRI’s performance under controlled conditions. Section 6 demonstrates its application to real ASKAP observations. Finally, Section 7 summarises the contributions and discusses potential directions for future work.
2 Hyperspectral RI imaging
@@
2.1 Hyperspectral radio emission model
In radio astronomy, continuum synchrotron emission typically varies smoothly with frequency, with flux density scaling according to a power-law (Thompson et al., 2017). This makes the power-law spectral model a practical choice for parametrizing the hyperspectral RI image cube. To account for sudden changes in the spectrum while still preserving overall spectral smoothness across a wide frequency range, a second-order term is introduced in the exponent (Rau & Cornwell, 2011). Consider a reference spectral frequency $\nu_{0}$ and the corresponding channel image $\overline{\bm{x}}_{0}\in\mathbb{R}^{N}$. At a given frequency $\nu_{l}$, the radio image $\overline{\bm{x}}_{l}\in\mathbb{R}^{N}$ can be modelled as
$\overline{\bm{x}}_{l}=\overline{\bm{x}}_{0}\left(\frac{\nu_{l}}{\nu_{0}}\right)^{-\bm{\alpha}+\bm{\beta}\log\left(\frac{\nu_{l}}{\nu_{0}}\right)}$ (1)
@@ 2.1 Hyperspectral radio emission model
This model is widely used for characterizing various radio sources, including galaxies, which yield different spectral indexes due to the different physical radiation mechanisms (De Gasperin et al., 2018). Furthermore, it can serve as a physics-informed prior in the reconstruction of the hyperspectral image cube, ensuring meaningful correlations across different spectral channels (Junklewitz et al., 2015; Ceccotti et al., 2023).
2.2 Hyperspectral Fourier sampling in RI
When probing the radio sky, RI measurements are acquired using arrays of antennas operating at multiple spectral channels. At a given observation frequency, each pair of antennas captures a noisy spatial Fourier intensity measurement, also known as a visibility. The corresponding Fourier mode is derived from the projection of the associated baseline onto the plane orthogonal to the line of sight, denoted by $\overline{\bm{b}}$ in the unit of meters (Thompson et al., 2017). The collection of visibilities from all baselines, accumulated over time, forms the so-called $uv$-coverage. For a given channel $l\in\{1,\dots,L\}$, the probed Fourier mode $\bm{b}_{l}$ is described in the unit of the observation wavelength such that

### v2

@@ Abstract
The next-generation radio-interferometric (RI) telescopes require imaging algorithms capable of forming high-resolution high-dynamic-range images from large data volumes spanning wide frequency bands. Recently, AIRI, a plug-and-play (PnP) approach taking the forward-backward algorithmic structure (FB), has demonstrated state-of-the-art performance in monochromatic RI imaging by alternating a data-fidelity step with a regularization step via learned denoisers. In this work, we introduce HyperAIRI, its hyperspectral extension, underpinned by learned hyperspectral denoisers enforcing a power-law spectral model. For each spectral channel, the HyperAIRI denoiser takes as input its current image estimate, alongside estimates of its two immediate neighboring channels and the spectral index map, and provides as output its associated denoised image. To ensure convergence of HyperAIRI, the denoisers are trained with a Jacobian regularization enforcing non-expansiveness. To accommodate varying dynamic ranges, we assemble a shelf of pre-trained denoisers, each tailored to a specific dynamic range. At each HyperAIRI iteration, the spectral channels of the target image cube are updated in parallel using dynamic-range-matched denoisers from the pre-trained shelf. The denoisers are also endowed with a spatial image faceting functionality, enabling scalability to varied image sizes. Additionally, we formally introduce Hyper-uSARA, a variant of the optimization-based algorithm HyperSARA, promoting joint sparsity across spectral channels via the $\ell_{2,1}$-norm, also adopting FB. We evaluate HyperAIRI’s performance on simulated and real observations. We showcase its superior performance compared to its optimization-based counterpart Hyper-uSARA, CLEAN’s hyperspectral variant in WSClean, and the monochromatic imaging algorithms AIRI and uSARA. HyperAIRI’s MATLAB implementation is available in the BASPLib code library.
[…]
facilities: VLA, ASKAP.software: PyTorch (A. Paszke et al., 2019), NUFFT (J. A. Fessler & B. P. Sutton, 2003), WSClean (A. Offringa & O. Smirnov, 2017), MeqTrees (J. E. Noordam & O. M. Smirnov, 2010), CASA (The CASA Team et al., 2022).
show]y.wiaux@hw.ac.uk
I Introduction
@@
Radio interferometry is a powerful technique in astronomy, providing high-resolution views of the sky by combining signals from an array of antennas. In particular, the upcoming Square Kilometre Array (SKA, A. M. M. Scaife, 2020) is designed to deliver unprecedented sensitivity and resolution, with phase 1 offering more than a tenfold improvement compared to its precursor instruments, MeerKAT (J. L. Jonas, 2009) and Australian SKA Pathfinder (ASKAP, A. Hotan et al., 2021). These telescopes observe across wide frequency bands by collecting Fourier measurements, enabling the study of astrophysical processes at both spatial and spectral scales.
@@
However, the associated inverse problem of hyperspectral image reconstruction is inherently ill-posed due to incomplete Fourier sampling, instrumental noise, and calibration errors. Monochromatic algorithms, which reconstruct spectral channels independently, achieve only limited recovery under such conditions. In contrast, joint-channel reconstruction exploits spectral correlations to regularize the ill-posed problem and enables high-precision recovery, particularly in high-dynamic-range (HDR) scenarios. Nevertheless, the sheer scale of hyperspectral data and the need for precise imaging significantly increase the computational demands, calling for efficient and scalable algorithms with joint-channel reconstruction capability.
@@
Most radio sources emit synchrotron radiation that can be described by a power-law spectral model (A. R. Thompson et al., 2017). Building on the celebrated CLEAN algorithm (J. Högbom, 1974), its first hyperspectral extension, Multi-Frequency (MF) CLEAN, models spectral variations via polynomial functions, and solves for the corresponding coefficient images during reconstruction (R. Sault & M. Wieringa, 1994). Further improvements are enabled by combining MF with Multi-Scale (MS) CLEAN (T. J. Cornwell, 2008), enhancing the recovery of extended emission (U. Rau & T. J. Cornwell, 2011; A. Offringa & O. Smirnov, 2017). However, low-order polynomials are often inadequate to capture complex spectral behavior (A. M. Scaife & G. H. Heald, 2012). To address this, E. Ceccotti et al. (2023) have injected a forced-spectrum fitting approach within CLEAN, enforcing a prior spectral knowledge directly during reconstruction. While computationally efficient, CLEAN-based algorithms still fall short in achieving precise imaging, i.e. high-resolution and high-sensitivity imaging. These limitations stem from their simplistic prior model: spatial resolution remains constrained by the instrument’s nominal resolution, and dynamic range is degraded by the addition of residuals in the final reconstruction.
@@
Bayesian inference methods, such as the variational Bayes approach underpinning the RESOLVE algorithm (H. Junklewitz et al., 2016), have demonstrated superior precision and robustness over CLEAN-based techniques (P. Arras et al., 2021; J. Roth et al., 2023). Assuming a first-order power-law spectral model, H. Junklewitz et al. (2015) proposed the joint estimation of a reference-frequency image and a spectral index map. Although these methods naturally provide uncertainty quantification functionality, their scalability and precision on large-scale observations remain to be demonstrated.
@@
Methods rooted in optimization theory have shown promise in astronomical imaging, offering precision imaging. Optimization provides a robust and versatile framework to address ill-posed inverse problems, typically formulated as minimization tasks solved via convergent iterative algorithmic structures. In this context, the reconstructed image is obtained as the minimizer of an objective function composed of a data fidelity function and a regularization function encoding sophisticated image models. Optimization offers a multitude of distributed algorithmic structures for parallelized reconstruction at scale. Among these methods, the SARA family of algorithms leverages sparsity-based priors, demonstrating superior imaging quality compared to CLEAN across diverse imaging scenarios (R. E. Carrillo et al., 2012; A. Onose et al., 2016, 2017; A. Repetti et al., 2017; J. Birdi et al., 2018; L. Pratley et al., 2018; A. Dabbech et al., 2018; M. Terris et al., 2023; A. G. Wilber et al., 2023a). Specifically, HyperSARA, promotes joint-channel average sparsity of the hyperspectral image cube in a redundant wavelet dictionary, and its low-rankness (A. Abdulaziz et al., 2016, 2019). HyperSARA takes the primal-dual forward-backward algorithmic structure (PD-FB, L. Condat, 2013), which enables full splitting of the functions involved, and parallel updates of the dual variables. To further improve scalability to large image sizes, P.-A. Thouvenin et al. (2023a) proposed its faceted variant operating on 3D spatial-spectral image facets. The algorithm has been validated on real observational data, demonstrating state-of-the-art reconstruction (P.-A. Thouvenin et al., 2023b). Yet, HyperSARA requires heavy cross-channel communications and could introduce significant overhead for extreme-scale problems. In this work, we introduce another simpler variant of HyperSARA, dubbed Hyper-uSARA, which adopts an unconstrained formulation of the minimization task, thereby eliminating the need for exact noise distribution knowledge. It also takes a simpler algorithmic structure, the FB (P. L. Combettes & J.-C. Pesquet, 2011), alternating data fidelity and regularization steps, while still allowing for full parallelization within each step.
@@
Recently, deep learning has gained increasing attention in RI imaging for its ability to produce reconstructions with rapid inference. Existing approaches focus on monochromatic imaging. These include traditional, end-to-end neural networks, either reconstructing images from back-projected data (L. Connor et al., 2022; C. Gheller & F. Vazza, 2022) or post-processing suboptimal outputs from other algorithms such as CLEAN (M. Terris et al., 2019). However, these methods often lack precision, robustness, and interpretability (I. J. Goodfellow et al., 2014; A. Nguyen et al., 2015; T. Pang et al., 2018). Conditioned denoising diffusion probabilistic models have also been explored for RI imaging (R. Wang et al., 2023; M. Drozdova et al., 2024). They face similar issues and remain in early development.
@@
Advanced architectures such as unrolling Deep Neural Networks (DNN) promote data consistency and possibly algorithmic stability by unfolding a fixed number of iterations of an optimization algorithm as key modules within the network (V. Monga et al., 2021). These networks are trained end-to-end while preserving an explicit link to optimization theory. However, existing unrolling schemes remain limited by their fixed iteration count, and lack formal convergence guarantees. Furthermore, embedding computationally-intensive measurement operators directly into the network can make training prohibitively expensive.
@@
In contrast, the Plug-and-Play (PnP) framework bridges optimization theory and deep learning (S. V. Venkatakrishnan et al., 2013; Y. Romano et al., 2017; E. T. Reehorst & P. Schniter, 2018) offering a more flexible and robust alternative. In this approach, proximal operators enforcing handcrafted regularization within an optimization algorithm are replaced by learned DNN denoisers. These denoisers can be agnostic to the measurement process, making them adaptable to diverse configurations in RI imaging. . AIRI has been validated on both simulated and real wide-field observational data (A. Dabbech et al., 2022; A. G. Wilber et al., 2023b). Thanks to their convolutional architecture and compact receptive fields, AIRI denoisers are typically trained on small image patches, and seamlessly applied to images of arbitrary size during inference. Additionally, they can be easily applied in parallel to image facets for scalability. These properties make AIRI well suited to large-scale precision imaging.
@@
Inspired by unrolled networks and PnP algorithms, A. Aghabiglou et al. (2024) have recently proposed R2D2, a novel imaging paradigm for radio interferometry. R2D2 is underpinned by a series of iteration-specific DNNs, each taking as input the current image estimate and associated data residual. The final reconstruction is formed by summing the DNNs’ output. In contrast to unrolled networks, R2D2 decouples data fidelity from the network architecture thereby avoiding the associated computational challenges. Conceptually, it can be interpreted as a learned version of CLEAN (A. Dabbech et al., 2024), substituting CLEAN’s minor cycles with iteration-specific DNNs. The novel paradigm has demonstrated superior imaging precision to the state of the art in field, namely AIRI and uSARA, both in simulation and on real data, only at a fraction of their computational cost. However, early incarnations of R2D2 have been so far trained in a telescope-specific setting, and for monochromatic intensity imaging at small scale. Extending it to large scale wideband imaging poses additional challenges that remain to be addressed.
@@
Building on the precision and scalability of AIRI, we propose HyperAIRI, a novel PnP algorithm tailored for high-precision hyperspectral RI imaging. HyperAIRI applies learned denoisers to each spectral channel in parallel with side information from neighboring channels. The reconstruction approach further enhances spectral coherence by incorporating a physical power-law model, with the spectral index map either estimated on the fly during reconstruction or provided as an input. This design ensures local communication at each iteration and requires global communication only during spectral index updates, making HyperAIRI highly scalable and efficient in distributed high-performance computing (HPC) environments. Similarly to AIRI, HyperAIRI takes FB, with convergence guaranteed by the non-expansiveness of the denoisers, enforced via a Jacobian regularization term in the training loss. We provide a shelf of pre-trained denoisers covering a wide dynamic-range, eliminating the need for problem-specific training. We demonstrate the superior performance of HyperAIRI, compared with state-of-the-art methods in the field, including the optimization-based Hyper-uSARA and joint-channel CLEAN, on simulated the Very Large Array (VLA, R. Perley et al., 2011) data and real ASKAP observations.
@@
The remainder of this paper is organized as follows. In Section II, we formalize the hyperspectral RI imaging problem. Section III reviews the hyperspectral SARA prior and formulates the Hyper-uSARA algorithm. Section IV presents the HyperAIRI algorithm, including the network architecture, training methodology, and overall algorithm structure. Section V reports experimental results on simulated datasets to assess HyperAIRI’s performance under controlled conditions. Section VI demonstrates its application to real ASKAP observations. Finally, Section VII summarizes the contributions and discusses potential directions for future work.
II Hyperspectral RI imaging
@@
II.1 Hyperspectral radio emission model
In radio astronomy, continuum synchrotron emission typically varies smoothly with frequency, with flux density scaling according to a power-law (A. R. Thompson et al., 2017). This makes the power-law spectral model a practical choice for parameterizing the hyperspectral RI image cube. To account for sudden changes in the spectrum while still preserving overall spectral smoothness across a wide frequency range, a second-order term is introduced in the exponent (U. Rau & T. J. Cornwell, 2011). Consider a reference spectral frequency $\nu_{0}$ and the corresponding channel image $\overline{\bm{x}}_{0}\in\mathbb{R}^{N}$. At a given frequency $\nu_{l}$, the radio image $\overline{\bm{x}}_{l}\in\mathbb{R}^{N}$ can be modeled as
$\overline{\bm{x}}_{l}=\overline{\bm{x}}_{0}\left({\nu_{l}}/{\nu_{0}}\right)^{-\bm{\alpha}+\bm{\beta}\log\left({\nu_{l}}/{\nu_{0}}\right)}$ (1)
@@ II.1 Hyperspectral radio emission model
This model is widely used for characterizing various radio sources, including galaxies, which yield different spectral indexes due to the different physical radiation mechanisms (F. De Gasperin et al., 2018). Furthermore, it can serve as a physics-informed prior in the reconstruction of the hyperspectral image cube, ensuring meaningful correlations across different spectral channels (H. Junklewitz et al., 2015; E. Ceccotti et al., 2023).
II.2 Hyperspectral Fourier sampling in RI
When probing the radio sky, RI measurements are acquired using arrays of antennas operating at multiple spectral channels. At a given observation frequency, each pair of antennas captures a noisy spatial Fourier intensity measurement, also known as a visibility. The corresponding Fourier mode is derived from the projection of the associated baseline onto the plane orthogonal to the line of sight, denoted by $\overline{\bm{b}}$ in the unit of meters (A. R. Thompson et al., 2017). The collection of visibilities from all baselines, accumulated over time, forms the so-called $uv$-coverage. For a given channel $l\in\{1,\dots,L\}$, the probed Fourier mode $\bm{b}_{l}$ is described in the unit of the observation wavelength such that